# Weak-Coupling Limit of the Lattice Nonlinear Schrödinger Integral Equation

### 10 Mar 2026 v1

@@ 4.3 Total density expansion
The full derivation is given in Appendix B. Numerically, fitting D(Q)QD(Q)-Q to $a\log Q+b$ for $Q\in[20,300]$ gives $a=0.1592\approx 1/(2\pi)=0.15915\ldots$
@@
with $b=-0.2173\pm 0.0005$ (estimated from variation of the fitting range; see Appendix B). The constant $b$ encodes contributions from both the edge boundary layers and the subleading corrections to the inner-region profile. Structurally, it should be expressible in terms of the Wiener–Hopf data of the symbol $\Sigma(p)=1-e^{-|p|}$—specifically, through the constant term $\beta_{\mathrm{FH}}$ in the asymptotic expansion of the Fredholm determinant $\log\mathcal{F}(Q)$ (see (114)), where the Fisher–Hartwig exponent is $\alpha_{\mathrm{FH}}=1/4$ corresponding to a simple zero $\Sigma(p)\sim|p|$ at the origin. For Fisher–Hartwig symbols of this type (a single zero of order $\alpha=1$), the constant $\beta_{\mathrm{FH}}$ is expressed in terms of the Barnes $G$-function evaluated at arguments related to the symbol’s value at the zero [13], typically taking the form $\beta_{\mathrm{FH}}=\log G(1+\alpha)\,G(1-\alpha)$ with $\alpha$ depending on the precise normalisation of the symbol. A search over combinations of elementary constants ($\gamma_{E}$, $\log 2$, $\log\pi$, $\pi$) does not match $b$ to better than $3\times 10^{-3}$, consistent with the expectation that $b$ involves $\beta_{\mathrm{FH}}$ and hence the Barnes $G$-function. Its analytical determination remains an open problem. At fixed particle density $D_{0}$, the Fermi boundary satisfies
@@ 5 Edge boundary layer
Let us define $s=Q-\xi\geq 0$ (distance from the right edge into the bulk). For $s=\mathcal{O}(1)$ and $Q\gg 1$, the contributions from the left edge at $\xi=-Q$ are exponentially small, and the integral equation reduces to a half-line convolution problem:
@@
where $g(Q-s)\approx 2/Q^{2}$ is essentially zero for $Q\gg 1$ and $K(x)=2/(1+x^{2})$ is the Lorentzian kernel. Taking the Fourier transform on $s\geq 0$ produces a Wiener–Hopf equation with the symbol
@@
The factorisation $\Sigma(p)=K_{+}(p)\,K_{-}(p)$, with $K_{+}$ ($K_{-}$) analytic and nonzero in the upper (lower) half-plane, can be obtained in closed form using the gamma-function identity $\Gamma(1+ix)\Gamma(1-ix)=\pi x/\sinh(\pi x)$. The result is (see Appendix C for the proof):
$K_{+}(z)$ $=\frac{(-iz)^{1/2}}{\Gamma\!\left(1-\frac{iz}{2\pi}\right)}\,\exp\!\left(-\frac{iz}{2\pi}\log(-iz)\right),\qquad\operatorname{Im}z>0\,$, (71)
$K_{-}(z)$ $=\frac{(iz)^{1/2}}{\Gamma\!\left(1+\frac{iz}{2\pi}\right)}\,\exp\!\left(\frac{iz}{2\pi}\log(iz)\right),\qquad\operatorname{Im}z<0\,.$ (72)
@@
Near the origin, $K_{+}(p)\sim(-ip)^{1/2}$ and $K_{-}(p)\sim(ip)^{1/2}$, so the linear zero of $\Sigma$ is distributed as a square-root singularity between the two factors. This square-root behaviour controls the edge profile: $\tilde{\rho}_{\mathrm{edge}}(s)\sim s^{1/2}$ as $s\to 0^{+}$, i.e. the density vanishes as a square root at the Fermi boundary.
It is convenient to define the regularised symbol $G(p)=(1-e^{-|p|})/|p|$, with factorisation $G=G_{+}\,G_{-}$ where $G_{\pm}=K_{\pm}/(\mp iz)^{1/2}$. Since $\Gamma(1)=1$ and the logarithmic exponentials tend to unity as $z\to 0$, one has $G_{+}(0)=G_{-}(0)=1$, so the regular part of the symbol is normalised to unity at the origin.
@@
$A_{\mathrm{WH}}=2\,,\qquad\log A_{\mathrm{WH}}=\log 2\,.$ (74)
This is the origin of the $\log 2$ in the constant $C=(\gamma_{E}+\log 2)/\pi$: the factor of $2$ comes from the ratio of the domain length ($2Q$) to the effective wavelength ($Q$), as dictated by the Wiener–Hopf data.
Normalising the edge profile by its value at a reference depth $s_{\mathrm{ref}}\gg 1$ in the outer region, solutions at different $Q$ collapse onto a universal curve $\Psi(s)$: $\tilde{\rho}(Q-s;\,Q)/\tilde{\rho}(Q-s_{\mathrm{ref}};\,Q)\to\Psi(s)$ as $Q\to\infty$, where $\Psi(s)\to 0$ as $s\to 0^{+}$ (the Fermi boundary) and $\Psi(s)\to 1$ as $s\to s_{\mathrm{ref}}$ (matching onto the outer region). The $s^{1/2}$ onset of $\Psi$ near $s=0$ reflects the square-root singularity of the Wiener–Hopf factors (71)–(72) at the origin. This universality reflects the fact that the edge equation (69) is independent of $Q$ to leading order; the collapse is illustrated in Figure 3.
Figure 3: Edge boundary layer at $\xi=Q$. (a) Raw edge profiles $\tilde{\rho}(Q-s;\,Q)$ versus the distance $s=Q-\xi$ from the Fermi boundary, for $Q=20$, $50$, $100$, $200$, and $400$. The overall scale decreases with $Q$ (reflecting the $Q$-dependent outer-region density), while the $s^{1/2}$ onset predicted by the Wiener–Hopf factorisation is visible at small $s$ (dashed guide). (b) Shape-collapsed profiles $\Psi(s)=\tilde{\rho}(Q{-}s;\,Q)/\tilde{\rho}(Q{-}s_{\mathrm{ref}};\,Q)$ with $s_{\mathrm{ref}}=10$. For $Q\geq 100$, the curves collapse onto a single universal profile, confirming the $Q$-independence of the edge boundary layer. The dashed line shows the $s^{1/2}$ behaviour near the Fermi boundary.
The three asymptotic regions developed in Sections 35—inner, outer, and edge—are summarised schematically in Figure 4. Having established the structure of the solution in each region and the matched asymptotic connections between them, we now turn to the analytical determination of the constant $C$.
@@ 6 Determination of $C$
where $\gamma_{E}=0.5772156649\ldots$ is the Euler–Mascheroni constant. We first present a heuristic derivation based on the digamma representation and mode counting, which motivates the result and provides physical intuition. We then give an analytical derivation using the Wiener–Hopf factorisation.
@@ 6.1 Mode-counting argument
On the full line ($Q=\infty$), the Fourier-space solution (34) gives the Bose–Einstein distribution. Expanding the Bose–Einstein kernel as a geometric series $1/(e^{p}-1)=\sum_{n=1}^{\infty}e^{-np}$ ($p>0$) and integrating term by term using
@@
This series diverges like $\sum 1/n$ for every $\xi$, reflecting the infrared divergence of the Bose–Einstein distribution.
We regularise by separating the divergent harmonic tail:
@@
$\tilde{\rho}_{\infty}(\xi)\;\sim\;\frac{1}{\pi}\bigl[\,H_{N}+S(\xi)\,\bigr]=\frac{1}{\pi}\bigl[\log N+\gamma_{E}-\operatorname{Re}\psi(1+i\xi)-\gamma_{E}\bigr]+\mathcal{O}(1/N)\,$, (84)
where $H_{N}=\sum_{n=1}^{N}1/n=\log N+\gamma_{E}+\mathcal{O}(1/N)$ is the $N$-th harmonic number and the sum is truncated at mode $n=N$ (the formal UV cutoff imposed by the finite domain).
The finite-domain equation on [Q,Q][-Q,Q] restricts the support of $\tilde{\rho}$ to a region of length $2Q$ in the rescaled variable $\xi$. The Lorentzian kernel $K(\xi)=2/(1+\xi^{2})$ has Fourier transform $\hat{K}(p)=2\pi e^{-|p|}$, which decays on a momentum scale of order 1. By the uncertainty principle for the truncated convolution operator, one expects the effective number of modes supported on [Q,Q][-Q,Q] to be
@@
assigning one mode per unit length—natural for a kernel with effective momentum bandwidth of order 1, though the specific prefactor requires justification beyond this heuristic. We note that this argument relies on two properties specific to the Lorentzian kernel: the momentum-space bandwidth is $\mathcal{O}(1)$ (since $\hat{K}(p)=2\pi e^{-|p|}$), and the regularised Wiener–Hopf factors satisfy $G_{\pm}(0)=1$ (see Section 6.2). For kernels with different bandwidth structure (e.g. a Gaussian kernel with $\hat{K}(p)=2\pi e^{-p^{2}}$), the effective number of modes would differ, and for kernels whose Wiener–Hopf normalisation gives $G(0)=c\neq 1$, an additional $\log c$ correction would appear. Evaluating (84) at $\xi=0$ with $N=2Q$ and using $S(0)=-\operatorname{Re}\psi(1)-\gamma_{E}=0$ (where $\psi(1)=-\gamma_{E}$):
$\tilde{\rho}(0;\,Q)$ $=\frac{1}{\pi}\bigl[H_{2Q}+S(0)\bigr]+\mathcal{O}\!\left(\frac{1}{Q}\right)$
$=\frac{1}{\pi}\bigl[\log(2Q)+\gamma_{E}\bigr]+\mathcal{O}\!\left(\frac{1}{Q}\right)$
@@
This suggests $C=(\gamma_{E}+\log 2)/\pi$, with the $\log 2$ arising from the factor of $2$ in $N_{\rm eff}=2Q$ (85). We now confirm this analytically.
@@ 6.2 Wiener–Hopf derivation of the constant
We now derive $C=(\gamma_{E}+\log 2)/\pi$ directly from the Wiener–Hopf factorisation of the symbol $\Sigma(p)=1-e^{-|p|}$, without appealing to the mode-counting heuristic. We proceed in three steps: (i) an exact spectral representation of $\tilde{\rho}(0;Q)$; (ii) identification of the spectral response function from the Wiener–Hopf data; (iii) evaluation of the resulting integral.
@@
Inserting into the eigenfunction expansion of the resolvent $(I-\mathcal{L}_{Q})^{-1}h$ at $\xi=0$ gives the exact, non-asymptotic identity
@@
Extend $\tilde{\rho}$ by zero outside [Q,Q][-Q,Q] and denote its Fourier transform $\hat{\rho}_{Q}(p)$. Since $\tilde{\rho}(0;Q)=(2\pi)^{-1}\!\int_{-\infty}^{\infty}\hat{\rho}_{Q}(p)\,dp$, it is natural to write
@@
where $\mathcal{R}(p;Q)$ is the spectral response function of the truncated interval, normalised so that $\mathcal{R}(p;Q)\to 1$ for each fixed $p>0$ as $Q\to\infty$, and $\mathcal{R}(0;Q)=0$ (reflecting the suppression of the zero mode by the finite domain). The full-line limit $\mathcal{R}\equiv 1$ recovers the divergent Bose–Einstein integral $\tilde{\rho}^{(\infty)}(0)=\pi^{-1}\!\int_{0}^{\infty}(e^{p}-1)^{-1}\,dp$.
To determine $\mathcal{R}$, we use the theory of truncated Wiener–Hopf operators on finite intervals [80, 13]. The key result, established in [80] (Chapter 4, specifically the finite-interval resolvent asymptotics for convolution equations with symbols vanishing at the origin), is that for the symmetric interval [Q,Q][-Q,Q] of length $L=2Q$, the truncated convolution equation has the structure of a paired Wiener–Hopf problem coupling the two edges at $\xi=\pm Q$. At low momentum $p\ll 1$, a Fourier mode propagates from the origin to one edge, is reflected with a coefficient determined by the Wiener–Hopf factors, propagates to the opposite edge, reflects again, and returns. The round-trip attenuation factor is $e^{-pL}=e^{-2pQ}$, multiplied by the product of the two reflection coefficients.
The reflection coefficient at each edge is controlled by the regularised Wiener–Hopf factors $G_{\pm}(z)=K_{\pm}(z)/(\mp iz)^{1/2}$. From the factorisation (71)–(72),
$G_{+}(z)=\frac{1}{\Gamma\!\bigl(1-\tfrac{iz}{2\pi}\bigr)}\;\exp\!\Bigl[-\frac{iz}{2\pi}\log(-iz)\Bigr],\qquad\operatorname{Im}z>0\,$, (91)
with $G_{-}(z)=\overline{G_{+}(\bar{z})}$ for real arguments. Since $\Gamma(1)=1$ and the logarithmic exponential tends to unity,
@@
$\log G_{+}(p)=\frac{ip}{2\pi}\bigl(\gamma_{E}-\log p\bigr)-\frac{p}{4}+\mathcal{O}(p^{2})\,$, (93)
so $|G_{+}(p)|^{2}=1-p/2+\mathcal{O}(p^{2})$ and the deviation from unity is $\mathcal{O}(p)$. The spectral response function is therefore
$\mathcal{R}(p;Q)=1-e^{-2pQ}\bigl[1+\mathcal{O}(p)\bigr]+\mathcal{O}(e^{-2\pi Q})\,$, (94)
where the $\mathcal{O}(e^{-2\pi Q})$ arises from the nearest complex zeros of $\Sigma(p)$ at $p=\pm 2\pi i$. The $\mathcal{O}(p)$ correction inside the brackets does not affect the leading asymptotics:
@@
since $\int_{0}^{\infty}p\,e^{-2pQ}/(e^{p}-1)\,dp=\mathcal{O}(1/Q^{2})+\mathcal{O}(1/Q)$ by dominated convergence (the integrand is bounded by $e^{-2pQ}$ near $p=0$ and decays exponentially for $p\gg 1$).
@@
Expanding $1/(e^{p}-1)=\sum_{n=1}^{\infty}e^{-np}$ and integrating term by term:
$\frac{1}{\pi}\int_{0}^{\infty}\frac{1-e^{-2pQ}}{e^{p}-1}\,dp=\frac{1}{\pi}\sum_{n=1}^{\infty}\left[\frac{1}{n}-\frac{1}{n+2Q}\right]=\frac{1}{\pi}\bigl[\psi(1+2Q)+\gamma_{E}\bigr],$ (97)
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is the digamma function and we used the classical series $\psi(1+z)+\gamma_{E}=\sum_{n=1}^{\infty}[1/n-1/(n+z)]$. Applying the asymptotic expansion $\psi(1+z)=\log z+1/(2z)+\mathcal{O}(1/z^{2})$ for $z=2Q\to\infty$:
$\tilde{\rho}(0;Q)=\frac{1}{\pi}\bigl[\log(2Q)+\gamma_{E}\bigr]+\mathcal{O}\!\left(\frac{1}{Q}\right)=\frac{\log Q}{\pi}+\frac{\gamma_{E}+\log 2}{\pi}+\mathcal{O}\!\left(\frac{1}{Q}\right).$ (98)
@@
The derivation identifies two independent contributions to $C$:
  • The Euler–Mascheroni constant $\gamma_{E}$ arises from the digamma asymptotics $\psi(1+z)=\log z+\gamma_{E}+\mathcal{O}(1/z)$, encoding the number-theoretic mismatch between the harmonic series and its logarithmic approximation. It is intrinsic to the $1/|p|$ singularity of the Bose–Einstein distribution and is independent of the boundary.
  • The $\log 2$ arises because the spectral response function (94) involves $e^{-2pQ}$, with the exponent $2Q$ equal to the full interval length $L=2Q$. This is dictated by the round-trip propagation across the domain [Q,Q][-Q,Q], with no additional multiplicative constant because $G_{\pm}(0)=1$. Had the regularised symbol satisfied $G(0)=c\neq 1$, the result would have been $C=(\gamma_{E}+\log 2+\log c)/\pi$.
The Wiener–Hopf normalisation $G_{\pm}(0)=1$ (92) is thus the structural reason that no further correction arises from the edge boundary layers beyond the geometric factor $L=2Q$. This confirms the heuristic mode-counting argument of Section 6.1 and explains why $N_{\rm eff}=2Q$ (85): the round-trip propagation factor is $e^{-2pQ}$ with unit reflection coefficients at each edge. Three-point Richardson extrapolation confirms $C=(\gamma_{E}+\log 2)/\pi$ to eight significant figures, as we now show.

### 22 Mar 2026 v2

@@ 4.3 Total density expansion
A companion identity controls the integrated density rather than the Lorentzian-weighted density. Integrating the digamma series
$\psi(1+z)+\gamma_{E}=\sum_{n=1}^{\infty}\left(\frac{1}{n}-\frac{1}{n+z}\right)$ (67)
with $z=i\xi$ term by term gives
$\int_{0}^{L}\operatorname{Re}\psi(1+i\xi)\,d\xi=\operatorname{Im}\log\Gamma(1+iL)\,$, (68)
since
$\int_{0}^{L}\frac{d\xi}{n+i\xi}=-i\log\frac{n+iL}{n}\,$, (69)
and resumming (by the Weierstrass product for $\log\Gamma$) we have
$\gamma_{E}L+\sum_{n=1}^{\infty}\left[\frac{L}{N}+i\log\left(\frac{n+iL}{n}\right)\right]=\log\Gamma(1+iL)\,.$ (70)
From (68) we can evaluate the integral of the profile function $\Phi(\xi)$ introduced in (50). From Stirling’s formula
$\operatorname{Im}\log\Gamma(1+iL)=L\log L-L+\frac{\pi}{4}+\mathcal{O}(1/L)\,$, (71)
we obtain
$\int_{0}^{\infty}\Phi(\xi)\,d\xi=-\frac{\pi}{4}\,.$ (72)
This identity will play a key role in the analysis of the constant $b$ in the total density expansion.
@@
The full derivation is given in Appendix B. Numerically, fitting D(Q)QD(Q)-Q to $a\log Q+b$ for $Q\in[20,300]$ gives $a=0.1592\approx 1/(2\pi)$.
@@
with $b=-0.2173\pm 0.0005$ (estimated from variation of the fitting range; see Appendix B). The structure of $b$ can be clarified using the profile integral (72). The additive composite of the inner and outer approximations is $\tilde{\rho}_{c}(\xi)=\frac{1}{2}-\Phi(\xi)/\pi$, where the first term is the Fermi-sea value (46) and the second accounts for the inner-region enhancement. Its integrated contribution to D(Q)QD(Q)-Q is
$2\int_{0}^{Q}\left[\tilde{\rho}_{c}(\xi)-\tfrac{1}{2}\right]d\xi=-\frac{2}{\pi}\int_{0}^{Q}\Phi(\xi)\,d\xi=\frac{1}{2}+\mathcal{O}(1/Q)\,$, (76)
by (72). Therefore the inner and outer regions together contribute precisely $1/2$ to D(Q)QD(Q)-Q, and the remaining content—the $(2\pi)^{-1}\log Q$ term and the constant bb—arises entirely from the edge boundary layers at $\xi=\pm Q$.
The computation of $b$ thus requires the subleading constant in the finite-interval Wiener–Hopf resolvent, which couples the logarithmic renormalisation of the effective domain length to the constant term through the Fisher–Hartwig correction to the Szegő asymptotics of the Fredholm determinant $\log\mathcal{F}(Q)$ (see (136)). For the symbol $\Sigma(p)=1-e^{-|p|}$ with its simple zero at the origin, the Fisher–Hartwig constant $\beta_{\mathrm{FH}}$ involves the Barnes $G$-function [13]. The analytical determination of $b$ remains an open problem. At fixed particle density $D_{0}$, the Fermi boundary satisfies
@@ 5 Edge boundary layer
Let us define $s:=Q-\xi\geq 0$, representing the distance from the right edge into the bulk. For $s=\mathcal{O}(1)$ and $Q\gg 1$, the contributions from the left edge at $\xi=-Q$ are exponentially small, and the integral equation reduces to a half-line convolution problem, i.e.,
@@
where $g(Q-s)\approx 2/Q^{2}$ is essentially zero for $Q\gg 1$. Taking the Fourier transform on $s\geq 0$ produces a Wiener–Hopf equation with the symbol
@@
The factorisation
$\Sigma(p)=K_{+}(p)\,K_{-}(p)\,$, (80)
with $K_{+}$ ($K_{-}$) analytic and nonzero in the upper (lower) half-plane, can be obtained in closed form, resulting in [^3] Ssee Appendix C for the proof.
$K_{+}(z)$ $=\frac{\sqrt{-iz}}{\Gamma\left(1-\frac{iz}{2\pi}\right)}\,\exp\left(-\frac{iz}{2\pi}\log(-iz)\right),\qquad\operatorname{Im}z>0\,$, (81)
$K_{-}(z)$ $=\frac{\sqrt{iz}}{\Gamma\left(1+\frac{iz}{2\pi}\right)}\,\exp\left(\frac{iz}{2\pi}\log(iz)\right),\qquad\operatorname{Im}z<0\,.$ (82)
@@
Near the origin, $K_{+}(p)\sim(-ip)^{1/2}$ and $K_{-}(p)\sim(ip)^{1/2}$, so the linear zero of $\Sigma(p)$ is distributed as a square-root singularity between the two factors. This square-root behaviour controls the edge profile: $\tilde{\rho}_{\mathrm{edge}}(s)\sim s^{1/2}$ as $s\to 0^{+}$, i.e., the density vanishes as a square root at the Fermi boundary.
It is convenient to define the regularised symbol
$G(p)=\frac{1-e^{-|p|}}{|p|}\,,\qquad G=G_{+}\,G_{-}\,,\qquad G_{\pm}=\frac{K_{\pm}}{\sqrt{\mp iz}}\,.$ (83)
Since the logarithmic exponentials tend to unity as $z\to 0$, one has $G_{+}(0)=G_{-}(0)=1$, so the regular part of the symbol is normalised to unity at the origin.
@@
$A_{\mathrm{WH}}=2\,.$ (85)
As we will discuss shortly, this is connected to the origin of the $\log 2$ in the constant (88). In other words, the factor of $2$ comes from the ratio of the domain length ($2Q$) to the effective wavelength ($Q$), as dictated by the Wiener–Hopf data.
Normalising the edge profile by its value at a reference depth $s_{\mathrm{ref}}\gg 1$ in the outer region, solutions at different $Q$ collapse onto a universal curve $\Psi(s)$:
$\frac{\tilde{\rho}(Q-s;\,Q)}{\tilde{\rho}(Q-s_{\mathrm{ref}};\,Q)}\to\Psi(s)$ (86)
as $Q\to\infty$, where $\Psi(s)\to 0$ as $s\to 0^{+}$ (the Fermi boundary) and $\Psi(s)\to 1$ as $s\to s_{\mathrm{ref}}$ (matching onto the outer region). The $\sqrt{s}$ onset of $\Psi$ near $s=0$ reflects the square-root singularity of the Wiener–Hopf factors (81)–(82) at the origin. This universality indicates the fact that the edge equation (78) is independent of $Q$ to leading order.
The three asymptotic regions developed in Sections 35—inner, outer, and edge—are summarised schematically in Figure 3.
@@ 6 Determination of $C$
where $\gamma_{E}$ is the Euler–Mascheroni constant. We first present a heuristic derivation based on the digamma representation and mode counting, which motivates the result and provides physical intuition. We then give an analytical derivation using the Wiener–Hopf factorisation.
@@ 6.1 Mode-counting argument
On the full line ($Q=\infty$), the Fourier-space solution (36) produces the Bose–Einstein distribution. Expanding it as a geometric series
$\frac{1}{e^{p}-1}=\sum_{n=1}^{\infty}e^{-np}\,,\qquad p>0\,$, (89)
and integrating term by term using
@@
The series diverges as $\sum 1/n$ for every $\xi$, reflecting the infrared divergence of the Bose–Einstein distribution. We can regularise it by separating the divergent harmonic tail:
@@
$\tilde{\rho}_{\infty}(\xi)\;\sim\;\frac{1}{\pi}\left[\,H_{N}+S(\xi)\,\right]=\frac{1}{\pi}\left[\log N+\gamma_{E}-\operatorname{Re}\psi(1+i\xi)-\gamma_{E}\right]+\mathcal{O}(1/N)\,$, (97)
where
$H_{N}=\sum_{n=1}^{N}\frac{1}{n}=\log N+\gamma_{E}+\mathcal{O}(1/N)$ (98)
is the $N$-th harmonic number and the sum is truncated at mode $n=N$ (the formal UV cutoff imposed by the finite domain).
The finite-domain equation on [Q,Q][-Q,Q] restricts the support of $\tilde{\rho}$ to a region of length $2Q$ in the rescaled variable $\xi$. The Lorentzian kernel (28) has Fourier transform
$\hat{K}(p)=2\pi e^{-|p|}\,.$ (99)
By the uncertainty principle for the truncated convolution operator, one expects the effective number of modes supported on [Q,Q][-Q,Q] to be
@@
assigning one mode per unit length—natural for a kernel with effective momentum bandwidth of order one. We note that this argument relies on two properties specific to the kernel: the momentum-space bandwidth is $\mathcal{O}(1)$ (see (99)), and the regularised Wiener–Hopf factors satisfy $G_{\pm}(0)=1$ (see Section 6.2). For kernels with different bandwidth structure (e.g. a Gaussian kernel with $\hat{K}(p)=2\pi e^{-p^{2}}$), the effective number of modes would differ, and for kernels whose Wiener–Hopf normalisation gives $G(0)=c\neq 1$, an additional $\log c$ correction would appear. Evaluating (97) at $\xi=0$ with $N=2Q$ and using $S(0)=-\operatorname{Re}\psi(1)-\gamma_{E}=0$ (where $\psi(1)=-\gamma_{E}$):
$\tilde{\rho}(0;\,Q)$ $=\frac{1}{\pi}\left[H_{2Q}+S(0)\right]+\mathcal{O}\left(\frac{1}{Q}\right)$
$=\frac{1}{\pi}\left[\log(2Q)+\gamma_{E}\right]+\mathcal{O}\left(\frac{1}{Q}\right)$
@@
This suggests (88), with the $\log 2$ arising from the factor of $2$ in $N_{\rm eff}=2Q$ (100). We now confirm this analytically.
@@ 6.2 Wiener–Hopf derivation of the constant
We now derive (88) directly from the Wiener–Hopf factorisation of the symbol (79). We proceed in three steps: (i) an exact spectral representation of $\tilde{\rho}(0;Q)$; (ii) identification of the spectral response function from the Wiener–Hopf data; (iii) evaluation of the resulting integral.
@@
Inserting into the eigenfunction expansion of the resolvent $(I-\mathcal{L}_{Q})^{-1}h$ at $\xi=0$ provides the identity
@@
Let us extend $\tilde{\rho}$ by zero outside [Q,Q][-Q,Q] and denote its Fourier transform $\hat{\rho}_{Q}(p)$. Since
$\tilde{\rho}(0;Q)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\hat{\rho}_{Q}(p)\,dp\,$, (105)
it is natural to write
@@
where $\mathcal{R}(p;Q)$ is the spectral response function of the truncated interval, normalised so that $\mathcal{R}(p;Q)\to 1$ for each fixed $p>0$ as $Q\to\infty$, and $\mathcal{R}(0;Q)=0$ (reflecting the suppression of the zero mode by the finite domain). The full-line limit $\mathcal{R}\equiv 1$ recovers the divergent Bose–Einstein integral
$\tilde{\rho}^{(\infty)}(0)=\frac{1}{\pi}\int_{0}^{\infty}\frac{dp}{e^{p}-1}\,.$ (107)
To determine $\mathcal{R}$, we use the theory of truncated Wiener–Hopf operators on finite intervals [80, 13]. The key result, established in [80] (Chapter 4, specifically the finite-interval resolvent asymptotics for convolution equations with symbols vanishing at the origin), is that for the symmetric interval [Q,Q][-Q,Q] of length $L=2Q$, the truncated convolution equation has the structure of a paired Wiener–Hopf problem coupling the two edges at $\xi=\pm Q$. At low momentum $p\ll 1$, a Fourier mode propagates from the origin to one edge, is reflected with a coefficient determined by the Wiener–Hopf factors, propagates to the opposite edge, reflects again, and returns. The round-trip attenuation factor is
$e^{-pL}=e^{-2pQ}\,$, (108)
multiplied by the product of the two reflection coefficients.
The reflection coefficient at each edge is controlled by the regularised Wiener–Hopf factors (83). From the factorisation (81)–(82),
$G_{+}(z)=\frac{1}{\Gamma\left(1-\tfrac{iz}{2\pi}\right)}\;\exp\left[-\frac{iz}{2\pi}\log(-iz)\right],\qquad\operatorname{Im}z>0\,$, (109)
with $G_{-}(z)=\overline{G_{+}(\bar{z})}$ for real arguments. As the logarithmic exponential tends to unity,
@@
$\log G_{+}(p)=\frac{ip}{2\pi}\left(\gamma_{E}-\log p\right)-\frac{p}{4}+\mathcal{O}(p^{2})\,$, (111)
so that
$|G_{+}(p)|^{2}=1-\frac{p}{2}+\mathcal{O}\left(p^{2}\right)\,$, (112)
and the deviation from unity is $\mathcal{O}(p)$. The spectral response function is therefore
$\mathcal{R}(p;Q)=1-e^{-2pQ}\left[1+\mathcal{O}(p)\right]+\mathcal{O}\left(e^{-2\pi Q}\right)\,$, (113)
where the $\mathcal{O}\left(e^{-2\pi Q}\right)$ arises from the nearest complex zeros of $\Sigma(p)$ at $p=\pm 2\pi i$. The $\mathcal{O}(p)$ correction inside the brackets does not affect the leading asymptotics:
@@
since
$\int_{0}^{\infty}\frac{p\,e^{-2pQ}}{e^{p}-1}\,dp=\mathcal{O}(1/Q^{2})+\mathcal{O}(1/Q)$ (115)
by dominated convergence (the integrand is bounded by $e^{-2pQ}$ near $p=0$ and decays exponentially for $p\gg 1$).
@@
Expanding via (89) and integrating term by term:
$\frac{1}{\pi}\int_{0}^{\infty}\frac{1-e^{-2pQ}}{e^{p}-1}\,dp=\frac{1}{\pi}\sum_{n=1}^{\infty}\left[\frac{1}{n}-\frac{1}{n+2Q}\right]=\frac{1}{\pi}\left[\psi(1+2Q)+\gamma_{E}\right],$ (117)
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is the digamma function and we used the classical series (67). Applying the asymptotic expansion
$\psi(1+z)=\log z+\frac{1}{2z}+\mathcal{O}(1/z^{2})$ (118)
for $z=2Q\to\infty$:
$\tilde{\rho}(0;Q)=\frac{1}{\pi}\left[\log(2Q)+\gamma_{E}\right]+\mathcal{O}\left(\frac{1}{Q}\right)=\frac{\log Q}{\pi}+\frac{\gamma_{E}+\log 2}{\pi}+\mathcal{O}\left(\frac{1}{Q}\right).$ (119)
@@
The derivation identifies two independent contributions to $C$: i) the Euler–Mascheroni constant $\gamma_{E}$ arises from the digamma asymptotics $\psi(1+z)=\log z+\gamma_{E}+\mathcal{O}(1/z)$, encoding the number-theoretic mismatch between the harmonic series and its logarithmic approximation. It is intrinsic to the $1/|p|$ singularity of the Bose–Einstein distribution and is independent of the boundary; ii) the $\log 2$ arises because the spectral response function (113) involves $e^{-2pQ}$, with the exponent $2Q$ equal to the full interval length $L=2Q$. This is dictated by the round-trip propagation across the domain [Q,Q][-Q,Q], with no additional multiplicative constant because $G_{\pm}(0)=1$. Had the regularised symbol satisfied $G(0)=c\neq 1$, the result would have been $C=(\gamma_{E}+\log 2+\log c)/\pi$. The Wiener–Hopf normalisation $G_{\pm}(0)=1$ (110) is thus the structural reason that no further correction arises from the edge boundary layers beyond the geometric factor $L=2Q$. This confirms the heuristic mode-counting argument of Section 6.1 and explains why $N_{\rm eff}=2Q$ (100): the round-trip propagation factor is $e^{-2pQ}$ with unit reflection coefficients at each edge.